# No Nowhere Constant Continuous Function Maps All Non-Normal Numbers to Normal Numbers

### v3

@@ 2 Proof of Main Results
are (relatively) open and let $\hat{U}_{x}$ further be the connected component of $U_{x}$ containing $x$. Then, the collection $(\hat{U}_{x})_{x\in I}$ forms an open cover of $I$. So, by compactness, there exists a finite cover $\hat{U}_{x_{1}},\ldots\hat{U}_{x_{n}}$ of $I$ with $x_{1}<x_{2}<\cdots<x_{n}$. Using again the assumption that $A$ is closed and nowhere dense, we can find open intervals $J_{x_{i},x_{i+1}}\subset(x_{i},x_{i+1}$ such that $J_{x_{i},x_{i+1}}\subset A^{c}$ for $i=1,\ldots,n-1$. Finally, we set $I_{x_{i}}:=\hat{U}_{x_{i}}\setminus\left(\bigcup_{j=1}^{n-1}J_{x_{j},x_{j+1}}\right)$ for $i=1,\ldots,n$. The intervals $I_{x_{i}},J_{x_{i},x_{i+1}}$ form a partition of $I$ and we define the function $g$ to be $g\equiv u_{x_{i}}$ on $I_{x_{i}}$ and the linear interpolation between $u_{x_{i}}$ and $u_{x_{i+1}}$ on $J_{x_{i},x_{i+1}}$. The function $g$ is clearly continuous and by construction $\|f-g\|<\varepsilon$ and $g(A\cap I)\cap A=\emptyset$. Since $\varepsilon>0$ is arbitrary, $Z$ is nowhere dense.
@@
Let us first demonstrate that the not simply $b$-normal numbers $\mathcal{N}_{b,s}^{c}$ contain a co-meager $G_{\delta}$-set $G_{b}$. Then, by Proposition 2.5, $G_{b}$ is super-dense and, thus, $\mathcal{N}_{b,s}^{c}$ is, too. As before, for $x\in{\mathbb{R}}$ we denote by $x_{i}\in\{0,\ldots,d-1\}$ the decimal places of $x$. We define
$G_{b}:=\bigcap_{n\geq 1}G_{n}$ (2.15)
The following proof follows the suggestion of the anonymous referee.
@@
Let $\varphi:\mathbb{R}\to\mathbb{R}$ be continuous and nowhere constant. By the intermediate value property, for every open interval $I$ there exists an open subinterval $I_{0}\subseteq I$ such that $\varphi(I_{0})$ contains an open set $J$ in its closure.
@@
$x\simeq y\Leftarrow x-y\in\mathbb{Q},$
which naturally decomposes $\mathbb{R}$} into equivalence classes $[x]$. We select exactly one representative of each equivalence class to form the set $V$. Of course, we made use of the axiom of choice in this construction. It is easy to see that we can assume $V$ to be dense. Indeed let $D\subset V$ be some infinitely countable set and $(I_{q_{1},q_{2}})_{q_{1}<q_{2}\in\mathbb{Q}}$ be the countable collection of open intervals with rational endpoints $q_{1},q_{2}$. After enumerating both countable sets, we replace each $d_{k}\in D$ by some $d_{k}+r_{k}\in I_{k}$ with $r_{k}\in\mathbb{Q}$. We fix such dense “Vitali” set $V$ and define $V_{q}:=\{x+q\,|\,x\in V\}$ for $q\in\mathbb{Q}$. By construction all $V_{q}$ are pairwise disjoint. We further introduce the following partition of the real line
@@
For the final assertion, we need a few set theoretic preliminaries. We denote as usual by $\aleph_{0}$ the cardinality of $\mathbb{N}$} and by the $\aleph_{1}$ the smallest uncountable cardinality, i.e., the cardinality of the set of all countable ordinal numbers. The cardinality of the the reals is given by $2^{\aleph_{0}}$ and under the continuum hypothesis $2^{\aleph_{0}}=\aleph_{1}.$ We further recall that any set $X$ can be well-ordered and for such a well-order $\prec$ we set $I_{\prec}(x)=\{t\in X\,|\,t\prec x\}$.
@@
We define $A:=\bigcup_{f\in F}A^{1}_{[}f]$ and due to the third property of $A_{f}^{1}$, the set $A$ is super-dense. Since $A^{c}$ contains all sets $A_{f}^{2}$, it is super-dense by the same reasoning. ∎
@@
for all $k\in S_{m}$. Here we used the function $\tau^{\prime}$ from the proof of Proposition 2.2. Our goal is to find $x^{(m+1)}\in I$ and $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\in T^{\prime}$ such that $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m$ and (2.16) holds true for $m+1$, too.
Denote by $U_{m}$ be the open set, for which $z\in U_{m}$ satisfies the condition (2.16) and $|z-x^{(m)}|<2^{-(m+2)}$. The standard argument involving the nowhere constancy and continuity of $\varphi_{\sigma(m+1)}$ and the density of $T^{\prime}$, we find an $x^{(m+1)}\in U_{m}$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}:=\varphi_{\sigma(m+1)}(x^{(m+1)})\in T^{\prime}$ with $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m.$
@@
The numbers $x^{(M)}$ form a Cauchy sequence with limit $x$. Since $x^{(1)}\in(1/4,3/4)$ and $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, we have $x\in(0,1)$. Since for any $k\in{\mathbb{N}}$ the sequence $y_{k}^{(m)}$ forms a subsequence of the function values $\varphi_{k}(x_{m})$, we have by continuity $y_{k}:=\varphi_{k}(x)\lim_{m\to\infty}y_{k}^{(m)}$. On the other hand,
$|y_{k}^{(m+1)}-y_{k}^{(m)}|<b^{-\tau^{\prime}\left(y_{k}^{(m)}\right)-1)},$
@@
    The set $\{\varphi_{k}^{\prime}(x)=0\}$ is finite and its size and elements are computable.
@@
we can be sure that there exists some $w\in[x^{(m)}-\delta_{m},x^{(m)}+\delta_{m}]$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}=\varphi_{\sigma(m+1),\varepsilon}(w)$. In contrast to the proof of Theorem 1.2, we cannot set $x^{(m+1)}=w$, but it is enough to find an $z$ such that (2.16) holds for $m+1$. We can find such an $x^{(m+1)}$ y a bisection procedure. ∎
Our last result deals with the non-normality for various bases.
Corollary 2.10.
$\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset$ (2.17)
@@
Appendix A Proof of Lemma 2.1
Proof of Lemma 2.1.

### v4

@@ 2 Proof of Main Results
are (relatively) open and let $\widehat{U}_{x}$ further be the connected component of $U_{x}$ containing $x$. Then, the collection $(\widehat{U}_{x})_{x\in I}$ forms an open cover of $I$. So, by compactness, there exists a finite cover $\widehat{U}_{x_{1}},\ldots\widehat{U}_{x_{n}}$ of $I$ with $x_{1}<x_{2}<\cdots<x_{n}$. Using again the assumption that $A$ is closed and nowhere dense, we can find open intervals $J_{x_{i},x_{i+1}}\subset(x_{i},x_{i+1})$ such that $J_{x_{i},x_{i+1}}\subset A^{c}$ for $i=1,\ldots,n-1$. Finally, we set $I_{x_{i}}:=\widehat{U}_{x_{i}}\setminus\left(\bigcup_{j=1}^{n-1}J_{x_{j},x_{j+1}}\right)$ for $i=1,\ldots,n$. The intervals $I_{x_{i}},J_{x_{i},x_{i+1}}$ form a partition of $I$ and we define the function $g$ to be $g\equiv u_{x_{i}}$ on $I_{x_{i}}$ and the linear interpolation between $u_{x_{i}}$ and $u_{x_{i+1}}$ on $J_{x_{i},x_{i+1}}$. The function $g$ is clearly continuous and by construction $\|f-g\|<\varepsilon$ and $g(A\cap I)\cap A=\emptyset$. Since $\varepsilon>0$ is arbitrary, $Z$ is nowhere dense.
@@
In fact, we can further strengthen the assertion of Proposition 2.1. Recall that a set $A\subset{\mathbb{R}}$ is said to be nowhere meager if $A\cap I$ is non-meager for every open interval $I\subset{\mathbb{R}}$.
Corollary 2.2.
Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is nowhere meager.
Proof.
We proceed via contradiction. Suppose $A\cap I$ is meager for an open interval $I$. Super-density is invariant under translation and stretching. Hence, we may assume without loss of generality that $B:=A\cap(0,1)$ is meager. We set
$\widehat{B}:=B\cup(1-B)=B\cup\{1-x\,|\,x\in B\}$
and further define the periodized
$\mathcal{B}:=\{x+n\,|\,x\in\widehat{B},n\in{\mathbb{Z}}\}.$
It is easy to see that $\mathcal{B}$ is meager, too. By Proposition 2.1, $\mathcal{B}$ is not super-dense and, thus, there exists a nowhere constant continuous function $\varphi:{\mathbb{R}}\to{\mathbb{R}}$ and an open interval $J\subset{\mathbb{R}}$ such that
$\varphi(J\cap\mathcal{B})\cap\mathcal{B}=\emptyset.$
Due to the periodicity of $\mathcal{B}$, we may assume that $J\subset(0,1)$. Consider the chainsaw-type function
$g(x):=\sum_{n\in 2{\mathbb{Z}}}|x-n|\mathbbm{1}_{[-1,1)}(x-n)$
which is exactly the periodic continuation of the absolute value from [1,1)[-1,1) to $\mathbb{R}$}. We observe that $g$ is continuous and $g\circ\varphi$ is still nowhere constant. Furthermore,
$(g\circ\varphi)(J\cap\mathcal{B})\cap\mathcal{B}=\emptyset.$
The reason is that we can write
$g(x)=\begin{cases}x-2n&\text{ if }x\in[2n,2n+1]\\ 1-(x-2n-1)&\text{ if }x\in[2n+1,2n+2]\end{cases}$
for $n\in{\mathbb{Z}}$ and the claim follows from the periodicity and reflection invariance of $\mathcal{B}$. We further shrink $J$ such that $\{0,1\}\notin(g\circ\varphi)(J)$. This is always possible due to continuity and the nowhere constancy of $g\circ\varphi$. Since $g\circ\varphi$ takes only values in $(0,1)$ on $J$ we obtain
$(g\circ\varphi)(J\cap A)\cap A=\emptyset.$
Here we used that $J\cap A\subset J\cap\mathcal{B}$ for $J\subset(0,1)$. This clearly shows that $A$ cannot be super-dense. ∎
@@
The following proof follows the suggestions of the anonymous referee.
@@
Let $\varphi:\mathbb{R}\to\mathbb{R}$ be continuous and nowhere constant. We claim that for every open interval $I$ there exists an open subinterval $I_{0}\subseteq I$ such that $\varphi(I_{0})$ contains an open set $J$ in its closure. Indeed, let us fix an open interval $I$ and we pick an $x\in I$. Since $I$ is open, we find an $\varepsilon>0$ such $(x-\varepsilon,x+\varepsilon)\subseteq I$. Since $\varphi$ is nowhere constant, we find an $y\in(x-\varepsilon,x+\varepsilon)$ such that $\varphi(x)\neq\varphi(y)$. We set $I_{0}=(x,y)$. We employ the intermediate value theorem to deduce that $\varphi(\overline{I_{0}})\supseteq[\varphi(x),\varphi(y)]$ and the latter clearly contains an open interval since $\varphi(x)\neq\varphi(y)$.
@@
In a next step, we combine our necessary and sufficient criteria from Corollary 2.2 and Proposition 2.3 to deduce the full characterization of super-dense with the Baire property, i.e. the first assertion of Theorem 1.3.
Corollary 2.4.
Suppose $A\subset{\mathbb{R}}$ has the Baire property. Then, $A$ is super-dense if and only if $A$ is co-meager.
Proof.
Suppose $A$ has the Baire property. We show both implications separately from each other.
Step 1: $A$ co-meager $\Rightarrow$ AA super-dense
Since $A^{c}$ is meager, there exists a countable family of nowhere dense sets $N_{k}$ such that $A^{c}=\bigcup_{k=1}^{\infty}N_{k}$. By definition, the closures $\overline{N}_{k}$ are nowhere dense, too. It follows that $\bigcup_{k=1}^{\infty}\overline{N}_{k}$ is a meager set and, hence,
$G:=\left(\bigcup_{k=1}^{\infty}\overline{N}_{k}\right)^{c}=\bigcap_{k=1}^{\infty}\overline{N}_{k}^{c}$
is a co-meager $G_{\delta}$-set which is contained in $A$. By Proposition 2.3, $G$ and thus $A$ is super-dense.
Step 2: $A$ super-dense $\Rightarrow$ AA co-meager
AA has the Baire property, so there exists an open set $U$ and a meager set $N$ such that $A=U\Delta N.$ We claim that $U$ is a dense open set. Indeed, if $U\cap I=\emptyset$ for an open interval, the Baire property implies that $A\cap I$ is meager. This contradicts the super-density of $A$ in view of Corollary 2.2. But this implies that $A^{c}\subset U^{c}\cup N$, where $U^{c}$ is closed with empty interior, i.e. $U^{c}$ is nowhere dense, and by definition $N$ is meager. In total, $A^{c}$ is meager. ∎
@@
$x\simeq y\iff x-y\in\mathbb{Q},$
which naturally decomposes $\mathbb{R}$} into equivalence classes $[x]$. We select exactly one representative of each equivalence class to form the set $V$. Of course, we made use of the axiom of choice in this construction. It is easy to see that we can assume $V$ to be dense. Indeed let $D\subset V$ be some infinitely countable set and $(I_{q^{(1)},q^{(2)}})_{q^{(1)}<q^{(2)}\in\mathbb{Q}}$ be the countable collection of open intervals with rational endpoints $q^{(1)},q^{(2)}$. Let $d_{1},d_{2},\ldots$ an enumeration of $D$ and similarly we fix an enumeration $I_{q_{k}^{(1)},q_{k}^{(2)}}$ of the rational open intervals. In a next step, we replace each $d_{k}\in D$ by some $d_{k}+r_{k}\in I_{q_{k}^{(1)},q_{k}^{(2)}}$ with $r_{k}\in\mathbb{Q}$. We fix such dense “Vitali” set $V$ and define $V_{q}:=\{x+q\,|\,x\in V\}$ for $q\in\mathbb{Q}$. By construction all $V_{q}$ are pairwise disjoint. We further introduce the following partition of the real line
@@
For the final assertion, we need a few set theoretic preliminaries. We denote as usual by $\aleph_{0}$ the cardinality of $\mathbb{N}$} and by the $\aleph_{1}$ the smallest uncountable cardinality, i.e., the cardinality of the set of all countable ordinal numbers. The cardinality of the reals is given by $2^{\aleph_{0}}$ and under the continuum hypothesis $2^{\aleph_{0}}=\aleph_{1}.$ We further recall that any set $X$ can be well-ordered which we denote by $\prec.$ For a fixed well-order $\prec$ and every $x\in X$, let $I_{\prec}(x)=\{t\in X\,|\,t\prec x\}$.
@@
We define $A:=\bigcup_{f\in F}A^{1}_{f}$ and due to the third property of $A_{f}^{1}$, the set $A$ is super-dense. Since $A^{c}$ contains all sets $A_{f}^{2}$, it is super-dense by the same reasoning. ∎
@@
Let us fix some notation. Given an integer base $b\geq 2$, we write $x=0.x_{1}x_{2}x_{3}\ldots$ for the $b$-ary expansion of $x\in[0,1)$. For $d\in\{0,1,\ldots,b-1\}$ and $M\in{\mathbb{N}}$, we set
$\Lambda_{b,d,M}(x):=\frac{1}{M}\sum_{j=1}^{M}\chi_{d}(x_{j})$ (2.3)
with the characteristic function $\chi_{d}$, that is, $\chi_{d}(d)=1$ and $\chi_{d}(k)=0$ for $k\neq d$. In other words, $\Lambda_{b,d,M}(x)$ is the density of the digit $d$ in the first $M$ digits of the $b$-ary expansion of $x$. For general $x\in{\mathbb{R}}$, we simply set $\Lambda_{b,d,M}(x):=\Lambda_{b,d,M}(|x|\,\mathrm{mod}\,1).$ We will be interested in the set $\mathcal{Z}_{b}$ of numbers with accumulations of zeros, that is,
Proposition 2.8 below gives a constructive demonstration for the super-density of $\mathcal{Z}_{b}$ and, thus, establishes the first part of Theorem 1.5.
Proposition 2.8.
Let $b\geq 2$ be an integer base, $I\subset{\mathbb{R}}$ a non-empty open interval, and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then, there is an iterative algorithm whose outputs converge to a real number $x$ with
$x\in\mathcal{Z}_{b}\,\text{ and }\,\varphi(x)\in\mathcal{Z}_{b}$ (2.5)
Proof.
of numbers with finite $b$-ary expansion and we set
$\tau(x):=\min\{L\in{\mathbb{N}}\,|\,x=0.x_{1}\ldots x_{L}\}$ (2.7)
for $x\in T$ and otherwise $\tau(x)=\infty$. Our construction requires a second set
where the inner union runs over all integers $K>\tau(x)^{2}+1$. A number $y\in T^{\prime}$ in the form $x+b^{-K}$ with $x\in T$ and $K>\tau(x)^{2}+1$ and this decomposition is in fact unique as $K$ is determined by the position of the last non-zero digit (which must be an isolated ’1’). For a $y=x+\frac{1}{b^{K}}\in T^{\prime}$, we similarly set $\tau^{\prime}(y):=K.$ The function $\tau^{\prime}(y)$ is well-defined by our previous observation. One easily sees that both, $T$ and $T^{\prime}$, are dense sets in $(0,1)$. The idea behind the definition of $T^{\prime}$ is that the zero digits of some $y=x+\frac{1}{b^{K}}\in T^{\prime}$ between the positions $\tau(x)$ and $K$ are preserved under small enough subtractions.
We now construct 4 sequences
Step 1: Choice of starting points and intervals
To construct the two other starting values, we recall that $\varphi$ is nowhere constant, and thus we find some $z\in I_{1,x}$ such that $\varphi(z)\neq\varphi(x_{1}^{(1)}).$ Employing the intermediate value theorem and the density of $T^{\prime}$, we in fact find some $x_{1}^{(2)}\in I_{1,x}$ such that $y_{1}^{(2)}:=\varphi(x_{1}^{(2)})\in T^{\prime}$. The digits $x_{1}^{(2)}$ at position $L_{x,1}$ to $L_{x,1}^{2}$-th are zero.
Step 2: Iterative procedure:
$I_{M+1,x}:=(x_{M+1}^{(1)},x_{M+1}^{(1)}+b^{-L^{2}_{x,M+1}-1})\cap I_{M,x}\neq\emptyset.$
[…]
$\Lambda_{b,0,L_{M,x}^{2}}(x)\geq 1-\frac{1}{L_{M,x}},$
@@
for all $k\in S_{m}$. Here we used the function $\tau^{\prime}$ from the proof of Proposition 2.8. Our goal is to find $x^{(m+1)}\in I$ and $y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\in T^{\prime}$ such that $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\right)>m$ and (2.9) holds true for $m+1$, too.
Denote by $U_{m}$ be the open set, for which $z\in U_{m}$ satisfies the condition (2.9) and $|z-x^{(m)}|<2^{-(m+2)}$. The standard argument involving the nowhere constancy and continuity of $\varphi_{\sigma(m+1)}$ and the density of $T^{\prime}$, we find an $x^{(m+1)}\in U_{m}$ such that $y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}:=\varphi_{\sigma(m+1)}(x^{(m+1)})\in T^{\prime}$ with $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\right)>m.$
@@
The numbers $x^{(M)}$ form a Cauchy sequence with limit $x$. Since $x^{(1)}\in(1/4,3/4)$ and $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, we have $x\in(0,1)$. Since for any $k\in{\mathbb{N}}$ the sequence $y_{k}^{(m)}$ forms a subsequence of the function values $\varphi_{k}(x_{m})$, we have by continuity $y_{k}:=\varphi_{k}(x)=\lim_{m\to\infty}y_{k}^{(m)}$. On the other hand,
$|y_{k}^{(m+1)}-y_{k}^{(m)}|<b^{-\tau^{\prime}\left(y_{k}^{(m)}\right)-1},$
@@
    The set $\{\varphi_{k}^{\prime}(x)=0\}$ is finite and its number of elements and their values are computable.
@@
we can be sure that there exists some $w\in[x^{(m)}-\delta_{m},x^{(m)}+\delta_{m}]$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}=\varphi_{\sigma(m+1),\varepsilon}(w)$. In contrast to the proof of Theorem 1.5, we cannot set $x^{(m+1)}=w$, but it is enough to find an $z$ such that (2.9) holds for $m+1$. We can find such an $x^{(m+1)}$ by a bisection procedure. ∎
2.3 Proof of Theorem 1.6
In this section, we construct an explicit function $\widehat{C}$ that contradicts the super-density of $\mathcal{N}$. Let us define
[…]
and recall that the set of $2$-simply normal numbers is given by
We set
$\mathcal{A}:=\mathcal{T}_{1}\cap\mathcal{N}_{s,2}.$ (2.10)
Proposition 2.10.
There is an explicit function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ such that
$\widehat{C}(\mathcal{A})\cap\mathcal{A}=\emptyset$ (2.11)
@@
Clearly, $C$ maps all numbers in $\mathcal{T}_{1}$ to values with finite binary expansion. That is,
$C(\mathcal{A}\cap(0,1))\cap\mathcal{A}=\emptyset.$ (2.12)
To this end, it is convenient to work with the symmetrized version $C_{s}$
which still is continuous and preserves the crucial property of the Cantor function that it maps all numbers with at least a ’1’ in the ternary expansion to a number with finite binary expansion. We now define the enriched Cantor function $\widehat{C}:[0,1]\to[0,1]$. To this end, we distinguish the cases, where $x$ has finitely many ’1’s and infinitely many. If $x$ has no ’1’, we set $\widehat{C}(x)=C_{s}(x).$ If $x$ has $m$ ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots w_{m}\sigma,$ where the finite words $w_{i}$ contain exactly an ’1’ as their last digit - we call such words one-stripped in this proof - and $\sigma$ is a word without any ’1’. Then, we set
$\widehat{C}(x):=\left(\sum_{r=1}^{m}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r})\right)+2^{-(m+1)\sum_{k=1}^{r}|w_{r}|}C_{s}(\sigma),$ (2.14)
$\widehat{C}(x):=\sum_{r=1}^{\infty}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r}).$ (2.15)
The function $\widehat{C}(x)$ is nowhere constant as $C(x)\neq C(y)$ for any $y$ with finite ternary expansion and any $x$ with infinitely many ’0’s and ’2’s.
Next, we verify that $\widehat{C}(x)$ is continuous. We consider the approximations $\widehat{C}_{k}(x),$ which roughly speaking stop after the $k$-th ’1’ (similarly to the original Cantor function which stops after the first ’1’). A rigorous construction goes as follows. Let $T_{k}:[0,1)\to[0,1)$ be the truncation map
[…]
where the $w_{i}$ are the unique one-stripped words from above with a single ’1’ as their last digit. We simply set $\widehat{C}_{k}(x):=\widehat{C}(T_{k}(x)).$ Let us show that $\widehat{C}_{k}(x)$ is continuous. We note first that $\widehat{C}_{1}(x)=C_{s}(x)$ and for each natural number $k$ the difference $\widehat{C}_{k+1}-\widehat{C}_{k}$ is nonzero only on closed intervals of the form
with one-stripped $w_{i}$ and on each such interval we have
$\widehat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)-\widehat{C}_{k}(0.w_{1}\ldots w_{k}\tau)=2^{-(k+1)\sum_{j=1}^{k}|w_{j}|}C_{s}(\tau)$
for any ternary word $\tau$. The continuity is evident as one inserts a damped copy of the original (symmetrized) Cantor function in each of the separated intervals $I_{w_{1},w_{2},\ldots,w_{k}}$ , where $\widehat{C}_{k}$ is constant. Moreover, by construction
$\sup_{x\in(0,1)}|\widehat{C}(x)-\widehat{C}_{k}(x)|\leq 2^{-k^{2}}$ (2.16)
so that $\widehat{C}_{k}$ converges uniformly to $\widehat{C}$. This implies that $\widehat{C}$ is continuous, too.
It remains to show that $\widehat{C}(\mathcal{T}_{1})\cap\mathcal{N}_{2,s}=\emptyset.$ To this end, let $x\in\mathcal{T}_{1}$ and we write for its ternary expansion $x=0.w_{1}w_{2}\cdots$ with one-stripped $w_{i}$. Let $\widehat{w_{i}}$ be the “Cantor image” of $w_{i}$ that is the binary string obtained by flipping each ’2’ in $w_{i}$ to a ’1’. Then, the binary expansion of $\widehat{C}(x)$ has the form $\widehat{C}(x)=\widehat{w_{1}}z_{1}\widehat{w_{2}}z_{2}\cdots$, where each $z_{r}$ is a string of zeros. Now, observe that after the last digit of $z_{m}$ the binary expansion of insertion of $\widehat{C}(x)$ has at most $\sum_{r=1}^{m}w_{r}$ ’1’s but at least $(m-1)\sum_{r=1}^{m}w_{r}$ zeros. That is there a sequence of integers $L_{m}$ s.t.
$\Lambda_{2,0,L_{m}}(\widehat{C}(x))\geq 1-\frac{1}{m}.$
This shows that $\widehat{C}$ has all desired properties on $[0,1]$. Since $\widehat{C}(0)=\widehat{C}(1)=0$, the periodic extension of $\widehat{C}$ to the whole real line is the desired function. ∎
In a next step, we want to show that the function $\widehat{C}$ is in fact computable. We follow the conventions from [23] where the reader may find more details. A name of a real number $x\in{\mathbb{R}}$ is a sequence of rational numbers $p_{n}$ with $|x-p_{n}|<2^{-n}$. A real function $f:I\to{\mathbb{R}}$ is computable if and only if there is a (Type-2) machine which transforms any name of $x\in I$ to a name of $f(x)$. Computability is a restrictive property as it requires the function to be continuous. We recall some key properties.
  1. 1.
    If $f,g$ are computable functions and $a$ is a computable real functions, then $af(x)+g(x),f(x)\cdot g(x),1/f(x)$ and $\min\{f(x),g(x)\}$ are also computable functions [23, Corollary 4.3.4].
  2. 2.
    The composition $f\circ g$ of two computable functions $f,g$ is computable (if well-defined) [23, Lemma 2.3.18].
  3. 3.
    Let $F:=\sum_{k=j}^{\infty}f_{j}$. Suppose that each $f_{j}$ is computable and that there exists a computable sequence of natural numbers $M_{k}$ such that $\sum_{j=M_{k}}^{\infty}\|f_{j}\|_{\infty}<2^{-k}$. Then, $F$ is computable [23, Theorem 4.3.8].
  4. 4.
    Suppose $f:[0,1]\to{\mathbb{R}}$ is computable and $f(0)=f(1)=0.$ The periodic continuation of $f$ is also computable. This follows for instance from [23, Lemma 4.3.5]
Theorem 1.6 follows from the following
Proposition 2.11.
The function $\widehat{C}$ from Proposition 2.10 is computable.
Proof.
Step 1: The Cantor function $C:[0,1]\to[0,1]$ is computable.
This fact is well-known, but hard to trace in the mathematical literature. We give a sketch of the proof. We construct a uniform Type–2 machine realizing $C$.
Let $\rho$ denote the standard Cauchy representation of $\mathbb{R}$ [23, Definition 4.1.5]. Thus a $\rho$–name of $x\in[0,1]$ is a sequence $(p_{n})_{n\in\mathbb{N}}$ of rationals satisfying $|p_{n}-x|\leq 2^{-n}$ for all $n$.
Define a partial operator
$M\colon\subseteq\mathbb{N}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$
as follows. On input a $\rho$–name $p=(p_{n})_{n}$ of $x$ and on input $k\in\mathbb{N}$, the machine $M$ proceeds uniformly:
  1. 1.
    Choose the least $n=n(k)$ such that $2^{-n}<3^{-(k+1)}$.
  2. 2.
    Using $p_{n}$, determine the unique integer $m\in\{0,\dots,3^{k}-1\}$ with
    $x\in\left[\frac{m}{3^{k}},\frac{m+1}{3^{k}}\right].$
    This uniquely fixes the first $k$ digits of the ternary expansion of $x$.
  3. 3.
    Replace each ternary digit 0 by 0 and each ternary digit $2$ by $1$. If a digit $1$ occurs, the resulting binary expansion is finite.
  4. 4.
    Let $q_{k}$ be the dyadic rational determined by the resulting $k$ binary digits and output $M(p)(k):=q_{k}$.
By construction,
$|q_{k}-C(x)|\leq 2^{-k}$
for all $k\in\mathbb{N}$. Hence $(q_{k})_{k}$ is a $\rho$–name of $C(x)$. The procedure is uniform and independent of the particular input $x$. Therefore $M$ is a realizer of $C$, i.e.
$\rho(M(p))=C(\rho(p))$
for all $\rho$–names $p$ of $x\in[0,1]$, and $C$ is $\rho$–computable.
Step 2: The function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ is computable.
This turns out to be largely a consequence of the above discussed properties of computability. Since $\widehat{C}$ is defined on $\mathbb{R}$} via a periodic extension, it is enough to show that $\widehat{C}:[0,1]\to{\mathbb{R}}$ is computable. Moreover, employing the functions $\widehat{C}_{k}$ from the proof of Proposition 2.10, we can write
$\widehat{C}(x)=\sum_{k=1}^{\infty}\widehat{C}_{k}(x)-\widehat{C}_{k-1}(x),$
where $\widehat{C}_{0}\equiv 0.$ In view of (2.16), it is enough to show that each difference $\widehat{C}_{k}-\widehat{C}_{k-1}$ is computable. For $k=1$, this follows from the representation $C_{s}(x)=\min\{1-C(x),C(x)\}$ and the computability of the Cantor function. For $k>1$, we use the series representation
$\widehat{C}_{k}(x)-\widehat{C}_{k-1}(x):=\sum_{w_{1},\ldots,w_{k-1}\text{ one-stripped}}\mathbbm{1}_{x=0.w_{1}\ldots w_{k-1}\tau}2^{-k\sum_{j=1}^{k-1}|w_{j}|}C_{s}(\tau).$
Note that each summand is computable as rescaled Cantor function. Due to the explicit damping, the full series is computable, too. This completes the proof. ∎
Appendix A Proof of $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$
The main purpose of this appendix is to present a proof of the following
Lemma A.1.
Let $b\geq 2$ be an integer. Then,
$\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0.$ (A.1)
The proof of Lemma A.1 relies on a standard combinatorial estimate and basic properties of Hausdorff measures. As it does not directly follow from Egglestone’s result on the Hausdorff-dimension of sets with specific asymptotic digit frequencies [10], we choose to sketch the complete argument.
Proof of Lemma A.1.